(4xA)+(6xB)=32
A+B=7

Answer by stanbon(48546)

4a+6b= 32
a+b=7
Use substitution to get:
a =7-b
Substitute to get:
4(7-b)+6b=32
28-4b+6b= 32
2b= 4
b=2
Substitute into a+b=7 to get
a = 5
So, a=5 and b=2
Cheers,
stan H.